OpenStudy (anonymous):

Still not getting it! The length of a rectangle is whatever the width is increased by 8 inches. The perimeter is 100 inches. find the length.

OpenStudy (aum):

Let the width be x inches. Length = x + 8 Perimeter = 2(length + width) = 2(x + 8 + x) = 2(2x+8) = 100 solve for x.

OpenStudy (anonymous):

x=21?

OpenStudy (anonymous):

21+8= the length?

OpenStudy (anonymous):

so, the length is the width plus 8 (w + 8) and the width is, well, the width (w). Now, the formula for the perimeter of a rectangle is P=2l + 2w. Because it has four sides, 2 of those are the length and 2 are the width. Following so far? So, if we were to set this up in a formula we would do it the same way.\[100=2(w+8) + 2w\] Because (w +8) is the measure of the length and w is the width. Now you just solve for the width. So, the width is equal to 21. BUT the question asks us for the LENGTH. So we just plug that into the length measure which is (w+8) which is 29. So the length equals 29. Does that make sense?

OpenStudy (anonymous):

thank you thank you! I wasn't sure if i was doing it correctly.